OpenStudy (anonymous):

A motor scooter purchased for $1,000 depreciates at an annual rate of 15%. Write an exponential function, and graph the function. Use the graph to predict when the value will fall below $100. 1. v(t) = 0.85(1000)t; the value will fall below $100 in about 18 yr. 2. v(t) = 1000(0.85)t; the value will fall below $100 in about 18 yr. 3. v(t) = 0.85(1000)t; the value will fall below $100 in about 14.2 yr. 4. v(t) = 1000(0.85)t; the value will fall below $100 in about 14.2 yr.

OpenStudy (anonymous):

but ecording to the equation the function must be exponential, how come the answer are given as a linear function?

OpenStudy (anonymous):

@JamenS

OpenStudy (anonymous):

@JamenS i suppose that t is an exponent. so let: v(t)=final prize of the scooter...t=time in years.....p=initial prize of the scooter . r= rate of depreciation per year so v(t)=p(1-r)^t.... does this help?

OpenStudy (phi):

Do you have a calculator?

OpenStudy (anonymous):

it loses 15% of its value, i.e. it retains 85% of its value use \(V=1000\times (.85)^t\)

OpenStudy (anonymous):

solve \(1000\times (.85)^t=100\) in two steps 1) divide both sides by \(1000\) to get \((.85)^t=0.1\) 2) use the change of base formula which says \(b^x=A\iff x=\frac{\ln(A)}{\ln(b)}\)

OpenStudy (anonymous):

oh third step is a calculator